OpenStudy (anonymous):

What is the 23rd term of the arithmetic sequence where a1 = 6 and a5 = 20?

OpenStudy (cwrw238):

common difference (d ) = (a5 - a1) / 4 = 20-6 / 4 = 3.5 23rd term = a1 + 22d = 6 + 22* 3.5

OpenStudy (anonymous):

is it 83?

OpenStudy (cwrw238):

6 + 77 = 83 yup

OpenStudy (anonymous):

thanks!